OpenStudy (gschibby):

Okay, matrices. How do I proceed when I have this expression: (1/3)*X*A+I2=A^T

OpenStudy (gschibby):

Written clean: \[\frac{ 1 }{ 3 }*X*A+I _{2}=A ^{T}\]

OpenStudy (gschibby):

And I need to find the matrice X

OpenStudy (anonymous):

Answer is AX/3 + 2i =A^t

OpenStudy (gschibby):

But I need to find X = ?

OpenStudy (gschibby):

And X*A is not equal to A*X ...